# imitation learning BEV bird’s-eye view

### v1

@@ III-C Probing Planning
    After each optimization step, determine if the planned trajectory results in a collision with any object at any future time step via an intersection check on the BEV bounding-boxes of all ego-object pairs. If no collision is found within $N_{\text{opt}}$ iterations, the optimization terminates unsuccessfully.
@@ IV Experiment Results
In this section, experimentally analyse Emperror in terms of (1) its effectiveness in faithfully imitating modern camera-based 3D object detectors and (2) illuminate its utility in probing the robustness of planning to such errors.
@@
Figure 3: PEM Qualitative Results. A comparison of a baseline PEM (middle) using an MLP with a Gaussian and Emperror (right) with DETR3D (left) as the target detector. Red boxes indicate ground truth objects, blue boxes the model predictions. While the baseline model samples implausible perception velocities and does not adapt to scene context, such as the parking area in Scene B, our approach closely mimics the target detector.
Target Detectors: To include a variety of error characteristics in our evaluation, we choose three different modern detectors: (1) DETR3D [2] based on sparse object queries, (2) StreamPETR [36] that utilizes temporal object queries that are propagated through time and (3) BEVFormer [29] which utilizes an intermediate temporal BEV-feature grid. All detectors have been trained for 48 epochs, utilizing a ResNet-101 [37] image backbone and the official implementations.
@@ IV-A Evaluation of Error Imitation Quality
In order to measure the performance of a PEM, we compare it in terms of the metrics defined above with the corresponding target detector. However, analyzing the mean precision and error values alone does not suffice, as it does not quantify how the errors evolve with decreasing detection confidence values. Hence, for all metrics we also compare the integral of absolute differences in the metric values over all recall intervals between the target detector and PEM, which we term Cumulative Absolute Difference Area (CD).
@@
Baselines: Inspired by the PEMs proposed in [16, 14] we design six baseline configurations that utilize a simple per-object feed-forward network to map the privileged ground-truth state to the target detection output. Besides the ground truth bounding boxes, the input state also contains a categorical object visibility level, as provided in the dataset [31]. Furthermore, we also construct a variant that additionally utilizes the same map encoder as our approach by concatenating the flattened map embedding to the per-object state projections. Note that these methods fail to explicitly model false positives [16, 14] and cannot capture duplicate detections, since each output depends on a corresponding ground truth object. The outputs of the baseline models are the confidence scores for each class, as well as the parameters of a probability distribution for all regression targets. We train the model by optimizing the negative log likelihood for a given target detector. For the network architecture, we use the same MLP as Emperror for the input state projection, followed by either a three layer MLP utilizing Layer normalization [39] and an ELU activation function [40] or a ResNet [37] as proposed in [14] without dropout. For the probability distribution, we either use a multivariate Gaussian as in [16] with a diagonal covariance matrix or a multivariate Student-T distribution as in [14] as the target distribution.
@@
Implementation Details: For Emperror, we use $N_{q}^{gt}=300$ queries for ground-truth, as well as $N_{q}^{ft}=128$ queries for false positives. Each query has a latent dimensionality of $d_{h}=256$ while the sampled latent code utilizes $d_{z}=32$. We choose $\beta=0.01$ for weighting $\mathcal{L}_{JS}$ after a warm-up period of three epochs, while setting $\alpha=0.5$ for an equal weighting of forward and backward KL. We filter detector targets and PEM outputs for a minimum confidence score $\max(\mathbf{b}^{cls})\geq 0.2$ to ignore low confidence predictions.
@@
Ablation Study: To verify our key design choices, the query initialization scheme and choice of divergence, we run experiments varying the number of false positive queries $N_{q}^{fp}$ and using the standard KL divergence instead of the skew-geometric JS divergence and report the results in Table I. Doubling the number of false-positive queries results in similar performance at a higher computational burden, while omitting them drastically degrades performance, highlighting their importance. When using the KL divergence instead of the skew-geometric JS Divergence we also see degraded performance. Additionally, we observe that, while the mean variance over latents per sample remains similar, the maximum variance can grow large when using the KL divergence, which provides an avenue for implausible attacks.
IV-B Adversarial Perception Errors
After demonstrating its advantages in faithfully imitating modern perception models in Section IV-A, we now apply Emperror to probe the robustness of the IL-based planner described Section III-B. To this end, we run our proposed adversarial optimization procedure on the scenes provided by the validation split, which is held-out for all involved models during training. To gain an understanding of its baseline performance, we first apply the planner to perception results obtained directly from the target detector. As additional validation, we also verify whether it performs similarly when operating on maximum likelihood samples obtained from Emperror (i.e. $\mathbf{z}_{n}=\boldsymbol{\mu}^{p}_{\theta}(\mathbf{S},\mathcal{M})\ \forall\ \mathbf{z}_{n}\in\mathbf{Z}$). Finally, we optimize for failure over varying latent space constraints $\kappa$.
Metrics: As we are interested in inducing unsafe behavior, we use the Collision Rate (CR) as the main metric for this experiment. It measures the percentage of scenes for which the planned trajectory collides with another object within the planning horizon $T_{\text{plan}}$. To better characterize the baseline performance of the planner, we also report the Average Displacement Error (ADE), which measures the average $l_{2}$-distance to the human expert trajectory within $T_{\text{plan}}$, as well as the Final Discplacement Error (FDE), which is similar to the ADE but considers only the final waypoint at $t=T_{\text{plan}}$.
@@ IV-B Adversarial Perception Errors
(a) Car
(b) Pedestrian
Figure 6: Post-Attack Error Characteristics. We show precision and mean regression errors for DETR3D and successful attacks (with $\kappa=3$).
@@
Results: In addition to the evaluation presented in Section IV-A, the quantitative results reported in Table II further indicate Emperror to be a capable perception proxy in terms of downstream planning. Compared to the actual inference results of the target detector, operating the planner on maximum likelihood proxy samples results in similar a CR, ADE and FDE. Furthermore, when allowing the adversarial optimization procedure to adjust the set of latent variables $\mathbf{Z}$, the baseline CR can be increased by up to $85$%, suggesting a vulnerability of the IL- and transformer-based planner to even plausible perturbations in its input. Since the planner is conditioned on driving straight-ahead in the majority of scenes due to the statistics of the dataset, the adversarial optimization most strongly affects the planned trajectory longitudinally, for example by inducing sudden dangerous braking or acceleration while in stationary traffic. This is visualized in Fig. 5. Additional examples can be found in the supplementary material.
Post-Attack Error Characteristics: Through likelihood regularization Eq. (6), the constraints placed on the search space Eq. (5) and use of the JS- instead of KL-Divergence Eq. (4), we design our method to be conservative such that it generates plausible attacks rather than extreme perturbations. This is evident from Fig. 6, where we plot the error characteristics for successful attacks (with $\kappa=3$): Even after adversarial optimization, perception results match the characteristics of the detector well, performing on-par or better, which highlights the conservative nature of our setup.
@@ V Conclusion
In this work, we presented Emperror, a novel generative perception error model (PEM) that mimics the outputs of a perception system given a ground-truth scene representation. We have demonstrated its utility as an evaluation tool by probing an imitation learning (IL)-based planner in terms of its robustness to noise in its inputs. Furthermore, we showed that Emperror more accurately captures the target detector’s error characteristics than PEMs used in previous work. However, there are remaining limitations opening avenues for improvement in future work. Firstly, we manually designed a cost function to induce a specific failure mode (i.e. collision). To enable the generation of critical perception errors inducing a wider range of planning failures, learning a general driving cost function from data, e.g. by means of inverse reinforcement learning [42], is an interesting direction. Secondly, extending Emperror to model latent features instead of explicit intermediate representations, such as 3D bounding boxes, would enable stress-testing modular end-to-end trainable architectures, which have recently gained popularity [34, 35].

### v2

@@ III-C Probing Planning
    After each step, determine if the planned trajectory results in a future collision with any object via an intersection check on the BEV bounding-boxes of all ego-object pairs. If no collision is found within $N_{\text{opt}}$ iterations, the optimization terminates unsuccessfully.
@@ IV Experiment Results
In this section, we experimentally analyze Emperror in terms of (1) its effectiveness in faithfully imitating modern camera-based 3D object detectors and (2) illuminate its utility in probing the robustness of planning to such errors.
@@
Figure 3: PEM Qualitative Results. Perception errors modeled by a baseline PEM (middle) using an MLP with a Gaussian and Emperror (right) compared to DETR3D detections (left). Red boxes indicate ground truth objects, blue boxes the model predictions. While the baseline model samples implausible perception velocities and does not adapt to scene context, such as the parking area in Scene B, our approach closely mimics the target detector.
Target Detectors: To include a variety of error characteristics in our evaluation, we choose three different modern detectors: (1) DETR3D [2] based on sparse object queries, (2) StreamPETR [3] that utilizes temporal object queries that are propagated through time and (3) BEVFormer [30] which utilizes an intermediate temporal BEV-feature grid. All detectors have been trained for 48 epochs, utilizing a ResNet-101 [38] backbone and the official implementations.
@@ IV-A Evaluation of Error Imitation Quality
Metrics: To measure the perception performance, we follow the official metric definitions of the nuScenes benchmark. These include the Average Precision (AP) as well as regression error metrics for true positives. The latter include the mean Average Translation Error (mATE), mean Average Orientation Error (mAOE) and mean Average Velocity Error (mAVE). For the exact metric definitions, we refer the reader to [32]. However, analyzing the mean precision and error values alone is insufficient, as it does not quantify how the errors evolve with decreasing detection confidence values. To this end, we compare the integral of absolute differences in metric values over all recall intervals between the target detector and PEM, which we term Cumulative Absolute Difference Area (CD).
@@
Baselines: We implement a simple baseline similar to the observation noise used in the Kalman filter, which models the perception error distribution as a scene-independent Gaussian distribution. To this end, we compute the per-class empirical Gaussian over 3D bounding box regression errors and class logits for detections with a ground-truth match on the PEM sub-split of the training dataset. Samples from this distribution are then applied to ground-truth bounding boxes to create noisy perception results. We additionally compute the per-class false-negative rate, at which we randomly drop detections. Since this baseline, which we term Static Gauss, simply models dataset-level statistics, it is incapable of capturing scene-dependent error patterns. Inspired by the PEMs proposed in [15, 13] we also design six object-conditioned baseline configurations that utilize a simple per-object feed-forward network to map the ground-truth state as well as a categorical visibility level [32] to a noisy detection output. Furthermore, we also construct a variant that additionally utilizes the same map encoder as our approach by concatenating the flattened map embedding to the per-object state projections. Note that these methods fail to explicitly model false positives [15, 13] and cannot capture error patterns that depend on other scene elements, such as duplicate detections. Unlike Static Gauss, they can, however, capture correlations between the ground-truth object state and detection errors on a per-object basis. These baselines output the confidence scores for each class, as well as the parameters of a probability distribution for all regression targets. We train the models by optimizing the negative log likelihood for a given target detector. We use the same MLP as Emperror for the input state projection, followed by either a three layer MLP utilizing Layer normalization [39] and an ELU activation function [40] or a ResNet [38] as proposed in [13] without dropout. For the probability distribution, we either use a multivariate Gaussian as in [15] with a diagonal covariance matrix or a multivariate Student-T distribution as in [13] as the target distribution.
@@
Implementation Details: For Emperror, we use $N_{q}^{gt}=300$ queries for ground-truth, as well as $N_{q}^{ft}=128$ queries for false positives. Each query has a latent dimensionality of $d_{h}=256$ while the sampled latent code utilizes $d_{z}=32$. This configuration is comparatively lightweight and achieves an inference throughput of roughly 190 scenes per second at batch size 64 on an Nvidia Titan Xp GPU. We choose $\beta=0.01$ for weighting $\mathcal{L}_{JS}$ after a warm-up period of three epochs, while setting $\alpha=0.5$ for an equal weighting of forward and backward KL. We filter detector targets and PEM outputs for a minimum confidence score $\max(\mathbf{b}^{cls})\geq 0.2$ to ignore low confidence predictions.
@@
Ablation Study: To verify our key design choices, the query initialization scheme and choice of divergence, we run experiments varying the number of false positive queries $N_{q}^{fp}$ and using the standard KL divergence instead of the skew-geometric JS divergence and report the results in Table I. Doubling the number of false-positive queries results in similar performance at a higher computational burden, while omitting them drastically degrades performance, highlighting their importance. When using the KL- instead of the $\text{JS}^{G_{\alpha}}$ divergence we also see degraded performance. Additionally, we observe that it can permit high latent variance, providing an avenue for implausible attacks.
@@
IV-B Adversarial Perception Errors
We now apply Emperror to probe the robustness of the IL-based planner described Section III-B. To this end, we run our proposed adversarial optimization procedure on scenes from the validation split, which is held-out for all involved models during training. To gain an understanding of its baseline performance, we first apply the planner to perception results obtained directly from the target detector. We also verify whether it performs similarly when operating on maximum likelihood samples obtained from Emperror (i.e. $\mathbf{z}_{n}=\boldsymbol{\mu}^{p}_{\theta,n}(\mathbf{S},\mathcal{M})\ \forall\ \mathbf{z}_{n}\in\mathbf{Z}$). Finally, we optimize for failure over varying latent space constraints $\kappa$.
Metrics: As we are interested in inducing unsafe behavior, we use the Collision Rate (CR) as the main metric for this experiment. It measures the percentage of scenes for which the planned trajectory collides with another object within the planning horizon $T_{\text{plan}}$. This is supplemented by the Average Displacement Error (ADE), which measures the average $l_{2}$-distance to the human expert trajectory within $T_{\text{plan}}$, as well as the Final Discplacement Error (FDE), which is similar to the ADE but considers only the final waypoint at $t=T_{\text{plan}}$.
Figure 6: Attack Plausibility. Successful attacks (with $\kappa=3$) show a similar or more accurate distribution of errors compared to DETR3D, highlighting the realism and conservative nature of our adversarial framework.
@@ IV-B Adversarial Perception Errors
Results: In addition to the evaluation presented in Section IV-A, the quantitative results reported in Table II further indicate Emperror to be a capable PEM in terms of downstream planning. Compared to the actual inference results of the target detector, operating the planner on maximum likelihood proxy samples results in similar a CR, ADE and FDE. Furthermore, when allowing the adversarial optimization procedure to adjust the set of latent variables $\mathbf{Z}$, the baseline CR can be increased by up to $85$%, suggesting a critical vulnerability of our prototypical planner to even plausible perturbations in its input. The adversarial optimization most strongly affects the planned trajectory longitudinally, for example by inducing sudden dangerous braking or acceleration while in stationary traffic. We hypothesize this is due to a bias of heading straight-ahead that is common in driving data. This is visualized in Fig. 5. Additional examples can be found in the supplementary material.
Attack Plausibility: Through likelihood regularization Eq. (6), search space constraints Eq. (5) and use of the JS- instead of KL-Divergence Eq. (4), we design our method to be conservative such that it generates plausible attacks rather than extreme perturbations. This is evident from Fig. 6, where we show boxplots of real-world detector errors and Emperror samples inducing successful attacks (with $\kappa=3$) on the same sensor logs: Even after adversarial optimization, Emperror produces results that either match the detector distribution or show lower errors, both in terms of the median and spread of the distribution, highlighting their plausibility.
@@ V Conclusion
We presented Emperror, a novel generative perception error model (PEM) that mimics the outputs of a perception system given a ground-truth scene representation. We have demonstrated its utility as an evaluation tool by probing an imitation learning (IL)-based planner in terms of its robustness to noise in its inputs. Furthermore, we showed that Emperror more accurately captures the target detector’s error characteristics than PEMs used in previous work. However, there are remaining limitations opening avenues for improvement in future work. Firstly, we manually designed a cost function to induce a specific failure mode (i.e. collision). To enable inducing a wider range of planning failures, learning a general driving cost function from data [42] is an interesting direction. Secondly, extending Emperror to model latent features instead of explicit intermediate representations, such as 3D bounding boxes, would enable stress-testing modular end-to-end trainable architectures, which have recently gained popularity [35, 36].